Session: Oct 2005 (session 3)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (An tao at Thu Oct 27 21:09:52 2005) Success :

Urumqi (Zhang hua at Thu Oct 27 21:26:33 2005) Success :

Torun (Kaz Borkowski at Fri Oct 28 06:23:54 2005) Success :
No known problems.

Onsala (Michael Lindqvist at Fri Oct 28 06:57:22 2005) Success :
Some RFI, otherwise no known problems.

Hartebeesthoek (Jonathan Quick at Fri Oct 28 07:35:39 2005) Minor failures :
Intermittent RFI, particularly in RCP

Medicina (Andrea Orlati at Fri Oct 28 08:01:48 2005) Success :
No known problems

Effelsberg (A. Kraus at Fri Oct 28 13:55:38 2005) Minor failures : FORM
First scans lost - formatter trouble. OK from UT1718 on.

Westerbork (Tony Foley at Mon Oct 31 08:39:01 2005) Minor failures :
Wb used in phased array mode. no known problems

Jodrell_Bank (Alastair Gunn at Tue Nov 1 12:20:46 2005) Success :
The Lovell Telescope was used. No known problems.

Cambridge (Alastair Gunn at Tue Nov 1 12:21:37 2005) Minor failures :
Mk5 kept dropping communications throughout experiment. Unit was restarted several times, resulting in data loss. Rack PSU was nearing failure which may have resulted in BBC unlocks, though FTP scans had fringes. Good luck with this one!

Robledo (Esther Moll at Wed Nov 2 14:33:31 2005) Minor failures : RFI POIN
VC#7 (281.99MHz) degraded by L-band known RFI. VC#5 (273.99MHz) was less affeted. Due to a encoder problem, the antenna was operated in computer mode instead of the usual mode. Pointing was checked before the observation. gps-fmout value at 16:14:31UT : 2005.300.16:14:31.04/gps-fmout/+7.676E-007 gps-fmout value at 19:48:21UT : 2005.300.19:48:21.04/gps-fmout/+7.756E-00

Arecibo (T. Ghosh/E.Momjian/C.J.Salter at Tue Nov 8 13:29:26 2005) Success :
Disk pack was sent to Jive on Nov-04